OpenStudy (anonymous):

derivative of cos(x/2)

OpenStudy (solomonzelman):

you will need to use the chain rule, for \(\large\color{black}{ \frac{\LARGE x}{\LARGE 2} }\) . I hope you know the derivative of just a, \(\large\color{black}{ \cos(x)}\), and know that \(\large\color{black}{ \frac{\LARGE x}{\LARGE 2} }\) is same as, \(\large\color{black}{ x\frac{\LARGE 1}{\LARGE 2} }\) .

OpenStudy (solomonzelman):

if you got any questions or concerns, you are free to ask....

OpenStudy (solomonzelman):

You don't need any more help? can you just in case tell me what you got?

OpenStudy (anonymous):

Thank You, The derivative of cosx is -sinx so I just need to apply the chain rule to that part of the equation to solve it.

OpenStudy (solomonzelman):

yes, you just need to apply the chain rule for the (x/2)

OpenStudy (solomonzelman):